OpenStudy (anonymous):

Solve for x. 6x - 3 + 12 = 5x + 24 - 4x A. x = -5 B. x = 2 C. x = 5 D. x = 3

pooja195 (pooja195):

Start by combining like terms

OpenStudy (anonymous):

Combining? You mean adding 5 on both sides or?

pooja195 (pooja195):

\[\huge~\rm~(6x)+(−3+12)=(5x−4x)+(24)\]

pooja195 (pooja195):

6x remains the same since nothing can be combined with it -3+12=?

OpenStudy (anonymous):

-3 + 12 = 9

pooja195 (pooja195):

Good \[\huge~\rm~6x+9=(5x−4x)+(24)\] Now the other side 5x-4x=?

OpenStudy (anonymous):

5x-4x=1x

pooja195 (pooja195):

Good and 24 remains the same \[\huge~\rm~6x+9=1x+24\] Ok whats the next step?

OpenStudy (anonymous):

add 6x + 9?

OpenStudy (anonymous):

Or am I diviiding?

pooja195 (pooja195):

You cant add those remember combine like terms We would do this: 

OpenStudy (anonymous):

Why would we be subtracting?

pooja195 (pooja195):

Because we want to get x by itself